OpenStudy (anonymous):

A bank received an initial deposit of $12,000, kept a percentage of this money in reserve based on a reserve rate of 6%, and loaned out the rest. The amount it loaned out eventually was all deposited back into the bank. If this cycle continued indefinitely, how much money eventually resulted from the initial deposit?

OpenStudy (jdoe0001):

I'm thinking the information is not enough to give a figure, or too ambiguous

OpenStudy (tkhunny):

12000*0.06= 720 12000 - 720 = 11280 11280*0.06 = 676.80 11280 - 676.80 = 10603.2 12000 11280 10603.2 etc. 12000 12000*0.94 12000*0.94*0.94 etc 12000(1 + 0.94 + 0.94^2 + ...) \(12000\cdot\dfrac{1}{1-0.94}\), or, we may wish to subtract the original 12000 from consideration.